### v5

@@ Abstract
This paper explores the exciting dynamics of international trade, focusing on the strategic balance between competition and cooperation in the global trade network. It highlights how competitive advantages, rather than comparative advantages, drive trade conflicts and deglobalization. By using the balance of power, the paper introduces a quantitative measure of competitiveness that serves as a common goal for all countries, alongside trade balance. It then examines how nations can boost their competitive strength and trade balance through globalization, protectionism, trade collaboration, or conflict. The study provides practical insights for policymakers on managing trade relationships, resolving conflicts, and determining the right level of globalization by analyzing trade data. The analysis is supported by a comparison between theory-driven quantitative evidence and historical events, using real-world trade data from 2000 to 2019.
@@
globalization, trade war, counterbalance equilibrium, national competitiveness, bargaining power, impossibility trilemma JEL Code: C71 , C78 , F11 , F13 , F15 , O24
Declarations of interests: I have nothing to declare.
Funding: This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Declaration of Generative AI and AI-assisted technologies in the writing process: During the preparation of this work, the author never used any Generative AI and AI-assisted technologies.
@@ 1 Introduction
Adam Smith’s concept of economies of scale and David Ricardo’s theory of comparative advantage suggest that global production would be maximized if all production factors could move freely. However, not all countries are equally prosperous; some succeed while others struggle. Even powerful economies have failed in the past, highlighting the need for fair profit sharing in international trade. Therefore, countries must compete for their share while cooperating to maximize global production. Both competition and cooperation are essential in international trade. This paper offers fair and strategic solutions for each nation within the trade network.
Analyzing trade wars and globalization presents several challenges. First, international trade involves both conflict and cooperation. While voluntary trade of goods and services results from bilateral collaboration, a country’s growing strength can weaken its trade partner’s position. Thus, a purely cooperative or non-cooperative study may not capture the complexity of trade relationships. The second challenge is creating a high-level description of the interactive trade system that focuses on competitive advantage, as exports and imports already reflect the comparative advantages of specific goods and services. The system is a directional network from which we aim to quantitatively formulate the economies’ utility and social welfare functions in this conflicting and cooperative context. Policy variables, such as imports and exports as percentages of production, should be linked to these economic objectives. Countries can then apply optimization methods to choose appropriate policies to achieve their economic goals. Finally, incorporating data into the analytics can explain past events and find optimal solutions for the near future. The solutions can identify the best trade partner, the right target in trade friction, a fair resolution to a conflict, or the optimal level of globalization.
We view the international trade system as a game on a network where countries act as nodes and trades as directional edges (e.g., Chaney, 2014; $\ddot{\mathrm{O}}$nder and Yilmazkuday, 2016). Trade volumes flow along the edges, with imports and exports moving in opposite directions. In the multilateral network, a unilateral action could achieve unpredictable or even opposite results. Trade wars or anti-globalization efforts are often framed to protect domestic industries or labor markets, reduce trade deficits, or overcome anti-dumping measures. These objectives, however, often result in more harm than gain (e.g., Chung, Lee, Osang, 2016; Read, 2005). Realistically, an attainable goal thrives on mutual understanding of the strengths and weaknesses of trade counterparts. For each country, the strengths and weaknesses lie in its domestic products and foreign imports and exports, forming a coalitional game. The Shapley value (1953) of the game provides a revealed cardinal preference or utility function for the individual country (Roth, 1977) and a power index in completing domestic production.
The common battlefield for competition and cooperation is a social welfare function, which weights the individual utility functions (e.g., Harsanyi, 1955; Hammond, 1987). These weights are chosen to balance the powers within the network (Hu and Shapley, 2003). The social welfare function measures the relative strengths of each country and serves as an objective in our analysis. It integrates both cooperation and competition within the global economic system, assessing the competitive advantages of countries and their long-term impacts on the entire system. This competitiveness generates additional welfare beyond economies of scale. The extra advantage, known as the Matthew effect, is another driving force for collaboration beyond comparative advantages.
De-globalization and trade wars represent two levels of economic conflict. The former involves a country’s withdrawal from the global economy, while the latter involves direct conflict between two trade partners. Both can negatively affect global economic activity. We apply a unified game-theoretic approach to formulate simple yet decisive rules to mitigate either type of conflict. Given that competitiveness has a constant sum in the game, the pursuit of competitiveness becomes the primary reason for initiating a trade war or anti-globalization action, while comparative advantages promote globalization and trade collaboration.
This study focuses on import and export activities, considering that geopolitical factors may already be incorporated into trade behaviors and volumes. Given the rationality of importers and exporters, trade data largely reflect political, territorial, ideological, cultural, war, national security, and other geopolitical considerations, alongside transaction and shipping costs, resource endowment, industrial distribution, and location advantages. From this perspective, a country’s primary objective in engaging in trade wars or de-globalization is to enhance its trade balance and competitiveness, as indicated by the data-driven social welfare function.
The objective of this paper is to offer strategic solutions for policymakers to enhance national competitive strength, rather than focusing solely on the competitiveness of individual firms. Government policymakers, rather than trade companies, are responsible for improving national competitiveness, as trade companies often exploit comparative advantages. A nation is not merely a collection of firms, and the interests of trade companies do not always align with those of policymakers, who must also consider the negative impacts of trade, such as the contraction of domestic labor markets and the relocation of industries overseas. Therefore, our study disregards production-based foundations (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014) and instead maximizes the social welfare function, emphasizing domestic production decision-making by global consumers. However, the measure of competitiveness is also a weighted average of comparative advantages, weighted by the preferences of consumers, whether affluent or impoverished.
@@
When a country initiates a trade war against a trade partner, it raises tariffs or imposes other barriers to enhance competitiveness. However, the outcome depends on the partner’s response. We find that a simple threshold for the counter-reaction — the ratio of their competitiveness — determines whether the country should engage in conflict or cooperation. Additionally, the reaction may benchmark their production ratio to avoid trade deficits. Utilizing the threshold and benchmark ratios, we derive Nash bargaining solutions (1950), generating consistent national bargaining power and providing methods to resolve bilateral trade frictions. A country with no specific targets to confront could choose between further globalization or protectionism. By making minor adjustments to its overall exports, we calibrate the effect on its competitiveness. The sign of the effect indicates whether the country should pursue further globalization. The analyses of trade wars and globalization also shed light on forming multilateral trade agreements.
The solutions are data-driven, relying solely on trade and production data. For a specific context with a set of fair assumptions, many researchers have studied bilateral or even multilateral negotiations (e.g., Bagwell, Staiger, and Yurukoglu, 2020; Schneider, 2005). There are also numerous debates over economic globalization in the literature, with various detailed viewpoints documented in Stiglitz (2002, 2017). Different perspectives, assumptions, and considerations can lead to opposing conclusions about globalization for a country. Our data-driven approach disregards arguments from both sides of the debates about labor markets, production chains, environmental issues, and national security. Trade policymakers may carefully weigh these considerations, and the trade data may already reflect them. Hence, a country could engage in further globalization or protectionism, depending on its particular situation at a specific time.
@@
We test our new theories by analyzing ComTrade data from the United Nations (2021) for the years 2000 through 2019, comparing theory-derived results with major economic events during the same period. Our findings suggest that protectionism could have been a more advantageous strategy for the USA and the United Kingdom (UK), but a less favorable option for China, Germany, Japan, and Russia, particularly if their objectives were centered on competitiveness. Data from 2017 indicate that trade wars might be a viable strategy for the USA to maintain its competitiveness and reduce its trade deficits, and we assess the impacts of the China–USA trade war in 2018 and 2019. However, we also observe significant adverse side effects on third parties. According to our theorem of impossibility trilemma, the USA must reduce imports if it aims to lower its national debt by generating trade surpluses while preserving its competitiveness.
@@
Our contribution to the trade literature lies in introducing a novel strategic and comprehensive framework for analyzing cooperation and competition within the global economic system. Our analyses eschew unnecessary assumptions and excessive mathematization, which can lead to misleading conclusions and policy implications (e.g., Romer, 2015). Unlike traditional studies, our approach does not rely on unknown parameters and formulates a quantitative measure of competitiveness for each country using the balance of power. This measure encompasses overall strength, including both direct and indirect influences. The study also derives national bargaining powers and offers strategic solutions for policymakers to manage trade relationships effectively.
@@
In the following exposition, we first introduce the competitiveness objective in Section 2. Section 3 formulates a necessary condition for a country to initiate a trade war or form a trade partnership to optimize its objectives. Section 4 derives Nash bargaining solutions to resolve bilateral conflicts or share profits. Section 5 discusses a similar condition for globalization or protectionism. In an empirical study of historical events, Section 6 examines the trade data from 2000 to 2019. Finally, Section 7 discusses several extensions and limitations of this framework. Our exposition is self-contained, and the proofs are in the Appendix.
@@ 2 A Network-based Social Welfare Function
This section introduces a novel type of competitiveness and social welfare function for countries within the trade network. Competitiveness is a relative measure derived from a constant-sum game. Absolute production and total trade volume may provide limited information regarding an economy’s competitive position. The game simulates the inflows and outflows within the network; without change, these steady flows reach a static equilibrium due to long-term accumulation. The equilibrium represents the competitiveness vector, which describes the utilities of the countries’ positions in the network game. We detail the steps toward the data-based social welfare function and its unique properties, which are unfamiliar in the trade literature.
@@
Domestic production is paramount for policymakers, whereas profit maximization is crucial for firms or producers. Production with zero or slightly negative profit may be insignificant to firm owners but can still be an option for the government to boost employment. For example, production maximization results in less unemployment, while labor may be resourced overseas to maximize shareholders’ profits. Thus, our objective function and strategic responses focus on the production decision-making process, leaving the flow of production factors and the distribution of profit to the firms. Furthermore, the government not only regulates profit-maximization behavior but also improves consumer welfare and redistributes the profits of domestic firms through taxation.
@@
For countries in the network, we describe their aggregate trade activities with a square matrix $P$. Assume there are $n$ countries, labeled as $1,2,\cdots,n$, and denote them collectively by the set $\mathcal{N}=\{1,2,\cdots,n\}$. For any $i,j\in\mathcal{N}$ with $i\neq j$, let $P_{ij}$ be the fraction of country $i$’s production that exports to country $j$. The exports include all services and goods, including intermediate goods; they may contain intermediate goods made by third parties or country $j$. Therefore, the production, denoted by $g_{i}$ for its value, consists of all types of exports as well as all domestically produced final products. Specifically, $g_{i}$ includes country $i$’s GDP and its imports from other countries, which are directly attributable to the production, such as raw materials, intermediate goods, and crude oil. It also includes exports of non-final products but excludes relevant imports, which are already included in $g_{i}$. Besides, $g_{i}$ should exclude exported final goods because they are already included in the GDP. Also, imported final goods which are not directly used in the domestic production should not be counted in $g_{i}$. We use $P_{ii}=1-\sum\limits_{j\neq i}P_{ij}\geq 0$ for the non-exporting fraction. Clearly, all $P_{ij}\geq 0$ for any $i,j\in\mathcal{N}$ and $\sum\limits_{j=1}^{n}P_{ij}=1$ for any $i\in\mathcal{N}$. We place all the fractions $P_{ij}$ into an $n\times n$ stochastic matrix $P=[P_{ij}]$.
@@
Each row of $P$ defines an individual utility function that determines a specific country’s production from a demand- or consumption-side perspective. For any $S\subseteq\mathcal{N}$, we define the set function $v_{i}:2^{\mathcal{N}}\to[0,1]$ by
@@
with the convention $v_{i}(\emptyset)=0$ for the empty set $\emptyset$. Thus, $v_{i}(\cdot)$ is a demand-driven production function for country $i$, and $v_{i}(S)$ is its value of production driven by the countries in $S$. Clearly, $(\mathcal{N},v_{i})$ defines a coalitional game, and its Shapley value (1953) is $(P_{i1},P_{i2},\cdots,P_{in})$, the $i$th row of $P$. Roth (1977) shows that the Shapley value is a von Neumann-Morgenstern utility function (1953), i.e., $P_{ij}$ measures an appraisal agent’s satisfaction when they play country $j$’s role in the game of $(\mathcal{N},v_{i})$. The agent is neutral with respect to ordinary and strategic risks detailed in Roth (1977). Thus, countries have cardinal preferences in the game where the inequality $P_{ij}>P_{ik}$ means country $j$ is preferred to $k$. Yet, their preferences change as $i$ varies in $\mathcal{N}$. In the matrix $P$, therefore, each row is an individual utility function, whereas column countries are the arguments or social states of the function. A social welfare function generally combines these individual utility functions into one, accounting for the diverse value judgments of countries in $\mathcal{N}$. It rates the set of countries in $\mathcal{N}$ as a whole and serves as a common platform for the countries to compete and collaborate.
@@ 2.1 Long-Run Influence in the Trade System
With the normalization (i.e., $v_{i}(\mathcal{N})=1$ and $v_{i}(\emptyset)=0$) and monotonicity of $v_{i}$, $P_{ij}$ bears much similarity to the Shapley-Shubik power index (1954), measuring country $j$’s chance of pivoting in forming the result $v_{i}(\mathcal{N})$. Indeed, $P_{ij}$ is a type of pro rata power when country $i$ makes its full production. The multilinear extension of $v_{i}$ is a function defined by:
$\tilde{v}_{i}(x_{1},x_{2},\cdots,x_{n})\overset{\mathrm{def}}{=\joinrel=}\sum\limits_{S\subseteq\mathcal{N}}\prod\limits_{j\in S}x_{j}\prod\limits_{k\not\in S}(1-x_{k})v_{i}(S)$
where $x_{j}\in[0,1]$ is the probability with which country $j$ participates in the consumption-driven production. Clearly, when restricted to $\{0,1\}^{n}$ or the corners of $[0,1]^{n}$, $\tilde{v}_{i}$ reduces to $v_{i}$. When the countries act independently, $\tilde{v}_{i}(x_{1},\cdots,x_{n})$ is country $i$’s expected production. Since the production process evolves from zero at the corner $(0,0,\cdots,0)$ to its completion at the corner $(1,1,\cdots,1)$, $x_{j}$ also represents $j$’s progress in completing its imports from $i$. Additionally, the partial derivative of $\tilde{v}_{i}$ with respect to $x_{j}$ measures country $j$’s instant influence or power to increase $\tilde{v}_{i}$ at $(x_{1},x_{2},\cdots,x_{n})$. Without any prior knowledge about the participation probabilities or progresses, we assume they advance evenly along the diagonal from $(0,\cdots,0)$ to $(1,\cdots,1)$. For example, if country $j$ completes 25% of its consumption, all other countries also finish their 25% at the same time. Consequently, the integral of $\tilde{v}_{i}$’s gradient $\bigtriangledown\tilde{v}_{i}$ alongside the diagonal quantifies each country’s overall influence or power that propels country $i$ to complete the production. According to Owen (1972), the integral equals the Shapley value for the coalitional game, i.e.,
@@
where $\top$” denotes vector or matrix transpose. Therefore, $P$ contains $n\times n$ quantitative powers, and the balance of power in $P$ refers to a state of stability.
@@
From a supply-side perspective, inputs to the production function in Eq. (1) can include any variables that affect the production and services of country $i$. However, it is impossible to compile a complete list of these variables. Their effects are also time-varying and country-specific; each product has a unique way of combining them. For example, inflows of financial assets, including remittances and foreign direct investments, are influential in making domestic products. These assets are primarily payments for past, current, or future exports. Trade involves exchanging goods and services; and monetary flow as the medium of exchange makes transactions possible in modern international trade. Nevertheless, payments alone cannot fully capture trade activities because exports and imports between two firms may be partially offset, resulting in net payments only. Zero payments may occur for intrafirm transactions of exports within a multinational company. Additionally, trade surpluses and deficits can result from manipulated currency depreciation and appreciation, respectively. A country can accumulate trade deficits if it is a highly desirable destination for foreign investment. For example, U.S. Treasury Bonds are lucrative and risk-free investment instruments after the Cold War, so foreign governments or companies must earn extra U.S. dollars from their trades to add to their foreign investment portfolios.
The direct influences in matrix $P$ automatically generate indirect ones in the closed global trade system. In a chain of production, even though the final product bears the mark of being made in one country, its components or parts may come from elsewhere. When China exports a smartphone to the USA, for example, the chip and design software of the phone could come from the USA, and the camera from Korea. Japan could manufacture the camera screen using US patents. To integrate these spillovers, we consider indirect and long-term impacts by the powers of $P$, such as $P^{2},P^{3},\cdots,P^{\infty}$. In general, $P^{t}$ contains the $t$-step aggregate impacts of all supply chains that crisscross the globalized world. For example, its value at the $i$th row and $j$th column, $\left(P^{t}\right)_{ij}$, accesses all the countries by $t-1$ times because
$\left(P^{t}\right)_{ij}=\sum_{k_{1}=1}^{n}\sum_{k_{2}=1}^{n}\cdots\sum_{k_{t-1}=1}^{n}P_{i,k_{1}}P_{k_{1},k_{2}}P_{k_{2},k_{3}}\cdots P_{k_{t-2},k_{t-1}}P_{k_{t-1},j}$ (2)
where $P_{k_{s},k_{s+1}}$ is the spillover of $k_{s}$’s production to $k_{s+1}$ for any $s=1,2,\cdots,t-1$. It is also $k_{s+1}$’s direct influence on $k_{s}$, which runs through all countries in $\mathcal{N}$.
@@
Under some general conditions (i.e., aperiodicity and irreducibility), the rows of $P^{\infty}$ are a constant row vector $\pi$, which exists uniquely. These conditions are obviously satisfied in the closed and well-connected trade system. Since $\pi_{j}=\lim\limits_{t\to\infty}\left(P^{t}\right)_{ij}$ regardless of the choice of $i$, $\pi_{j}$ measures country $j$’s long-run influence on the whole trade system. The direct influences in $P$ pass through interlinked global value chains and there are millions of them. Consequently, $P^{t}$ converges only if $t$ is large enough. The existence and uniqueness of $\pi$ can be found in the theory of Markov chains (e.g., Karlin and Taylor, 2012). The limit of $P^{t}$ provides a computational method to find $\pi$, and its speed of convergence is linear.

### v6

@@ Abstract
This paper investigates the evolving dynamics of international trade, emphasizing the strategic interplay between competition and cooperation within the global trade network. It argues that competitive advantages — rather than traditional comparative advantages — are the primary drivers of trade conflicts and deglobalization. Drawing on the concept of the balance of power, the paper introduces a quantitative measure of competitiveness, which complements the trade balance as a long-term policy objective. It further explores how countries can enhance competitiveness and trade balance through globalization, protectionism, collaboration, or trade frictions. Using real-world trade data from 2000 to 2019, our empirical study finds parallels between historical developments and quantitative evidence based on this new theory, offering actionable insights to policymakers for managing trade relations, mitigating conflicts, and calibrating the optimal level of globalization.
@@
globalization, trade war, counterbalance equilibrium, competitive advantage, bargaining power, impossibility trilemma JEL Code: C71 , C78 , F11 , F13 , F15 , O24
journal: Economic Analysis and Policy
@@ 1 Introduction
Adam Smith’s concept of economies of scale and David Ricardo’s theory of comparative advantage suggest that global production could be maximized if all factors of production moved freely. Yet, prosperity remains uneven: some nations flourish while others falter. Even historically dominant economies have faced setbacks, underscoring the importance of fair competition. To navigate this landscape, nations must strike a balance between competition and cooperation — securing their fair share while contributing to global output. Both dynamics are essential.
This paper proposes strategic and fair solutions for countries engaged in international trade. We identify three key challenges:
  1. 1.
    The Dual Nature of Trade: International trade embodies both conflict and cooperation. While voluntary exchange reflects collaboration, a country’s growing strength can undermine its partners. Models that are purely cooperative or non-cooperative fail to capture this complexity.
  2. 2.
    System-Level Representation: A framework that emphasizes competitiveness is crucial, as trade flows already reflect comparative advantages. We conceptualize international trade as a directional network and aim to formulate utility functions that integrate both conflict and cooperation. Policy variables — such as imports and exports as shares of production — should align with these objectives, enabling countries to optimize their strategies.
  3. 3.
    Data-Driven Insights: Real-world data can illuminate past events and guide optimal strategies for the near future — such as selecting trade partners, resolving disputes, and determining the appropriate degree of globalization.
We model international trade as a network game, where countries are nodes and trade flows are directional edges (e.g., Chaney, 2014; $\ddot{\mathrm{O}}$nder and Yilmazkuday, 2016). In this multilateral setting, unilateral actions often lead to unpredictable or counterproductive outcomes. Trade wars and anti-globalization measures — though intended to protect domestic industries or reduce deficits — frequently cause more harm than good (e.g., Chung et al., 2016; Read, 2005).
A realistic approach requires mutual understanding of each partner’s strengths and weaknesses, as reflected in domestic production and trade flows. These flows, along with domestic consumption, form a coalitional game for each nation. The Shapley value (1953) of the game provides a revealed preference or individual utility function (Roth, 1977) and serves as a power index in completing domestic production.
The commonly shared utility function — referred to as the social welfare function — aggregates individual utility functions (e.g., Harsanyi, 1955; Hammond, 1987) and balances power within the network (Hu and Shapley, 2003). It evaluates national competitiveness and captures the dual dynamics of cooperation and competition, thereby providing a battleground for both across nations.
Competitiveness generates additional welfare beyond economies of scale — a phenomenon often described as the Matthew effect — and offers further incentive for collaboration, complementing traditional comparative advantages. While comparative advantage refers to producing goods or services at a lower opportunity cost, thereby promoting specialization and inclusive growth, competitive advantage emphasizes strategic management aimed at achieving national dominance, fostering exclusive growth.
We adopt a unified game-theoretic framework to analyze two distinct forms of economic conflict: deglobalization and trade wars. Deglobalization refers to a country’s withdrawal from the global economy, whereas trade wars involve direct confrontation between trading partners. Both phenomena can significantly disrupt global economic activity. This paper proposes clear and practical strategies to mitigate these conflicts. In our framework, national competitiveness is modeled as a constant-sum game, suggesting that its pursuit often fuels trade wars and anti-globalization measures. In contract, comparative advantage promotes globalization and trade cooperation.
Our approach centers on trade flows, recognizing that geopolitical factors — such as political, territorial, ideological, cultural, and security considerations — are inherently embedded in trade data. These data also reflect transaction costs, resource endowments, industrial structures, and locational advantages. From this perspective, a country’s primary objective in trade wars or deglobalization is to improve its trade balance and competitiveness.
The goal of this paper is to offer strategic guidance for policymakers to enhance national competitiveness, rather than focusing on firm-level performance. Unlike trade firms, which primarily exploit comparative advantages, nations are not mere aggregations of firms. Policymakers must also address the adverse effects of trade, such as labor market disruptions and manufacturing offshoring. Accordingly, this study moves beyond production-based foundations (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014) and instead seeks to maximize a social welfare function that emphasizes domestic production decisions shaped by global demand. This function incorporates comparative advantages, weighted by consumer preferences across nations, products, and income levels.
@@
When a country initiates a trade war, it typically raises tariffs or imposes barriers to boost competitiveness. The outcome depends on the partner’s response. We find that a simple threshold — the ratio of competitiveness — determines whether conflict or cooperation is optimal. Additionally, responses may benchmark production ratios to avoid trade deficits. Using these threshold and benchmark ratios, we derive Nash bargaining solutions (1950) for fair responses. These solutions establish consistent measures of national bargaining power and offer practical methods for resolving bilateral trade frictions.
For countries without explicit confrontation targets, the strategic choice lies between globalization and protectionism. By adjusting overall exports, we can assess the impact on competitiveness; the sign of this effect indicates whether deeper globalization is beneficial. Additionally, our analyses of trade wars, globalization, and bargaining power all help form and reform multilateral trade agreements.
The proposed solutions are data-driven, relying exclusively on trade and production data. While prior research has examined bilateral or multilateral negotiations under specific assumptions (e.g., Bagwell et al., 2020; Schneider, 2005), and globalization has been extensively debated (e.g., Stiglitz, 2002 and 2017), these discussions often yield opposing conclusions depending on assumptions and priorities. In contrast, our data-driven approach sidesteps normative debates on labor markets, supply chains, environmental concerns, and national security — factors that trade data may already implicitly capture. Ultimately, whether a country embraces further globalization or protectionism depends on its unique circumstances at a given point of time.
@@
Using UN ComTrade data (2021), we test our theories against major economic events from 2000 to 2019. Results suggest that protectionism could have benefited the United States and the United Kingdom, but not China, Germany, Japan, or Russia, when competitiveness was the primary goal. Data from 2017 indicate that trade wars might have helped the USA maintain competitiveness and reduce deficits. We also assess the China–USA trade war in 2018 and 2019, noting significant spillover effects on third-party economies. Our impossibility trilemma shows that the USA should reduce imports if it aims to lower national debt by generating trade surpluses while preserving competitiveness.
@@
This study contributes to the trade literature by introducing a strategic, comprehensive framework for analyzing cooperation and competition in global trade. Unlike traditional studies, it avoids unknown parameters, unnecessary assumptions, and excessive mathematization, which can lead to misleading conclusions and policy implications (e.g., Romer, 2015).
@@
The remainder of this paper is organized as follows. We introduce the revealed social welfare function in Section 2. Section 3 formulates conditions for a country to initiate a trade war or form a trade partnership to optimize its competitiveness. Section 4 derives Nash bargaining solutions for resolving bilateral conflicts and sharing profits. Section 5 discusses conditions for choosing globalization or protectionism. Using trade data from 2000 to 2019, Section 6 presents an empirical analysis of historical events. Finally, Section 7 concludes with policy implications, extensions, and limitations of the framework. All proofs are provided in the Appendix, and the exposition is self-contained.
@@ 2 A Network-based Social Welfare Function
This section introduces a novel concept of competitiveness, which is a social welfare function for countries within the global trade network. Competitiveness is defined as a relative measure derived from a constant-sum game. Unlike absolute production or total trade volume, which provide limited insight into an economy’s competitive position, this measure captures the strategic dynamics of trade.
The game employs inflows and outflows within the network. In the absence of changes, these flows converge to a static equilibrium over time, reflecting long-term accumulation. This equilibrium yields a competitiveness vector, which represents the utilities of countries’ positions in the network game. We outline the steps for constructing this data-driven social welfare function and highlight its unique properties, which are largely unfamiliar in the trade literature.
@@
For policymakers, domestic production is a primary concern, whereas firms prioritize profit maximization. Production that yields zero or slightly negative profit may be unattractive to firms but can still be a viable policy option for governments seeking to boost employment. For example, maximizing production can reduce unemployment, while firms might prefer outsourcing labor overseas to maximize shareholder returns. Accordingly, our objective function and strategic responses focus on production decision-making, leaving the allocation of production factors and profit distribution to firms. Governments, however, play a dual role: regulating profit-maximization behavior and enhancing consumer welfare, while redistributing firm profits through taxation.
@@
To model trade relationships, we represent aggregate trade activities using a square matrix $P$. Assume there are $n$ countries, labeled $1,2,\cdots,n$, collectively denoted by the set $\mathcal{N}=\{1,2,\cdots,n\}$. For any $i,j\in\mathcal{N}$ with $i\neq j$, let $P_{ij}$ denote the fraction of country $i$’s production exported to country $j$. Exports include all goods and services, including intermediate goods, which may incorporate components from third countries or from country $j$ itself. The production value of country $i$, denoted $g_{i}$, consists of all types of exports, all domestically produced final products, and imports directly attributable to production (e.g., raw materials, intermediate goods, crude oil). However, $g_{i}$ excludes imported final goods not used in domestic production. We define $P_{ii}=1-\sum\limits_{j\neq i}P_{ij}\geq 0$ as the non-exporting fraction. Clearly, $P_{ij}\geq 0$ for all $i,j\in\mathcal{N}$, and for each $i\in\mathcal{N}$, $\sum\limits_{j=1}^{n}P_{ij}=1$. All fractions $P_{ij}$ form an $n\times n$ stochastic matrix $P=[P_{ij}]$.
@@
Each row of the matrix $P$ defines an individual utility function that breaks down a specific country’s production from either a demand-sided or consumption-side perspective. For any subset $S\subseteq\mathcal{N}$, we define the set function $v_{i}:2^{\mathcal{N}}\to[0,1]$ by
@@
with the convention $v_{i}(\emptyset)=0$ for the empty set $\emptyset$. Thus, $v_{i}(\cdot)$ is a demand-driven production function for country $i$, and $v_{i}(S)$ represents the value of production driven by the countries in $S$.
Clearly, $(\mathcal{N},v_{i})$ defines a coalitional game, and its Shapley value (1953) is given by the $i$th row of $P$, i.e., $(P_{i1},P_{i2},\cdots,P_{in})$. Roth (1977) demonstrates that the Shapley value is a von Neumann-Morgenstern utility function (1953). In this context, $P_{ij}$ measures an appraisal agent’s satisfaction when playing the role of country $j$ in the game of $(\mathcal{N},v_{i})$, assuming neutrality toward both ordinary and strategic risks (Roth, 1977). In the matrix $P$, therefore, rows represent individual utility functions of countries and columns represent the arguments or social states of these functions.
Consequently, countries exhibit cardinal preferences in the game: $P_{ij}>P_{ik}$ implies that country $j$ is preferred to $k$. However, these preferences vary with $i\in\mathcal{N}$. A social welfare function aggregates these individual utility functions into a single function, incorporating the diverse value judgments of countries in $\mathcal{N}$. It evaluates the set of countries as a whole and serves as a common platform for competition and cooperation.
@@ 2.1 Long-Run Influence in the Trade System
With the normalization conditions $v_{i}(\mathcal{N})=1$ and $v_{i}(\emptyset)=0$, and the monotonicity of $v_{i}$, the measure $P_{ij}$ closely resembles the Shapley-Shubik power index (1954), which quantifies country $j$’s pivotal role in determining the outcome $v_{i}(\mathcal{N})$. In essence, $P_{ij}$ represents a form of pro rata power when country $i$ realizes its full production.
The multilinear extension of $v_{i}$ is defined as:
$\tilde{v}_{i}(x_{1},x_{2},\cdots,x_{n})\overset{\mathrm{def}}{=\joinrel=}\sum\limits_{S\subseteq\mathcal{N}}\prod\limits_{j\in S}x_{j}\prod\limits_{k\not\in S}(1-x_{k})v_{i}(S).$
Here, $x_{j}\in[0,1]$ denotes the probability that country $j$ participates in the consumption-driven production. When restricted to the corners of the unit cube $0,1]^{n}$, $\tilde{v}_{i}$ simplifies to the original function $v_{i}$. Under the assumption of independent actions, $\tilde{v}_{i}(x_{1},\cdots,x_{n})$ represents the expected production of country $i$.
As production evolves from zero at the origin $(0,0,\cdots,0)$ to full completion at $(1,1,\cdots,1)$, each $x_{j}$ also reflects country $j$’s progress in completing its imports from country $i$. The partial derivative of $\tilde{v}_{i}$ with respect to $x_{j}$ captures country $j$’s instant influence or power to increase $\tilde{v}_{i}$ at a given point.
In the absence of prior knowledge about participation probabilities or progress, we assume uniform advancement along the diagonal from $(0,\cdots,0)$ to $(1,\cdots,1)$. For instance, if country $j$ completes 25% of its consumption, all other countries are assumed to have completed 25% as well. Consequently, the integral of the gradient of $\tilde{v}_{i}$, denoted $\bigtriangledown\tilde{v}_{i}$, along this diagonal quantifies each country’s overall influence in driving country $i$’s production to completion. According to Owen (1972), this integral equals the Shapley value for the coalitional game:
@@
where $\top$” denotes vector or matrix transpose. Therefore, this matrix $P$ contains $n\times n$ quantitative measures of power, and the balance of power in $P$ indicates a stable state in the trade system.
@@
From a supply-side perspective, the inputs to the production function in Eq. (1) may include any variables that influence the production and services of country $i$. However, compiling a comprehensive list of these variables is impractical. Their effects vary over time and across countries, and each product integrates them in unique ways. For example, inflows of financial assets — such as remittances and foreign direct investments — play a significant role in domestic production. These assets often represent payments for past, current, or future exports. Trade involves the exchange of goods and services; and monetary flows facilitate these transactions in modern international trade. Yet, payments alone do not fully capture trade activities. Transactions between two firms may be partially offset, resulting in net payments only. Intra-firm trade within multinational corporations may involve zero payments. Moreover, trade surpluses and deficits can be influenced by currency manipulation — depreciation and appreciation, respectively.
A country may accumulate trade deficits if it is a highly attractive destination for foreign investment. For instance, U.S. Treasury Bonds became lucrative and low-risk investment instruments after the Cold War. As a result, foreign governments or companies often seek to earn additional U.S. dollars through trade to expand their investment portfolios.
In a closed global trade system, direct influences captured in matrix $P$ naturally give rise to indirect ones . Within a production chain, even if the final product is labeled as originating from a single country, its components often come from multiple sources. For instance, when China exports a smartphone to the USA, the chip and design software may be sourced from the USA, the camera from Korea, and the camera screen manufactured in Japan using US patents.
To account for these spillovers, we consider both indirect and long-term impacts through successive powers of $P$, such as $P^{2},P^{3},\cdots,P^{\infty}$. In general, $P^{t}$ represents the aggregate $t$-step impacts across all supply chains in the globalized economy. Specifically, the entry at the $i$th row and $j$th column of $P^{t}$, denoted $\left(P^{t}\right)_{ij}$, captures all paths of length $t$ from country $i$ to country $j$:
$\left(P^{t}\right)_{ij}=\sum_{k_{1}=1}^{n}\sum_{k_{2}=1}^{n}\cdots\sum_{k_{t-1}=1}^{n}P_{i,k_{1}}P_{k_{1},k_{2}}P_{k_{2},k_{3}}\cdots P_{k_{t-2},k_{t-1}}P_{k_{t-1},j}.$ (2)
Here, each term $P_{k_{s},k_{s+1}}$ represents the spillover from country $k_{s}$ to $k_{s+1}$, or equivalently, the direct influence of $k_{s+1}$ on $k_{s}$, for $s=1,2,\cdots,t-1$. These interactions span all countries in the set $\mathcal{N}$.
@@
Under general conditions — specifically, aperiodicity and irreducibility — the rows of $P^{\infty}$ converge to a unique constant row vector $\pi$. These conditions are naturally satisfied in our trade system. The element $\pi_{j}$ is defined as: $\pi_{j}=\lim\limits_{t\to\infty}\left(P^{t}\right)_{ij}$ and is independent of the choice of $i$, thereby measuring country $j$’s long-run influence on the entire trade system. The existence and uniqueness of $\pi$ are well-established in the theory of Markov chains (e.g., Karlin and Taylor, 2012). Given the vast number of interlinked global value chains, convergence of $P^{t}$ requires a sufficiently large $t$. Moreover, the limit of $P^{t}$ offers a practical computational approach to determine $\pi$.